OpenStudy (anonymous):

Which of the following systems of equations has no solution? A. 18x − 6y = 14; 3y = 9x − 7 B. 7x − 4 = 3y; 6y − 3 = 14x C. 4x + y = 10; 2y = 8x − 2 D. -5y = x − 6; 3x − 15y = 9

OpenStudy (anonymous):

The way to go about doing this is to just manually solve every single option. It is tedious and simple at the same time.... A. 18x-6y = 14 3y = 9x - 7 => (times 2) 6y = 18x - 14 18x - (18x -14) = 14 14= 14 (all solutions of x and y) B. 7x − 4 = 3y => (times 2) 14x- 8 = 6y 6y − 3 = 14x (14x- 8) -3 = 14x 14x - 8 -3 = 14x 14x - 11 = 14x -11 = 0 (no solutions) Answer is B.

OpenStudy (anonymous):

0